HYDROGEN-BOUND COMPLEXES OF TROPOLONE: GATEWAYS FOR THE INTERROGATION OF MULTIPLE PROTON-TRANSFER EVENTS

\underline{DEACON J. NEMCHICK}, KATHRYN CHEW, JOHN E. WOLFF, and PATRICK H. VACCARO, Department of Chemistry, Yale University, P.O. Box 208017, New Haven, CT 06520-8107 USA.

Tropolone (TrOH) serves as a model system for the study of coherent proton-transfer processes, where a potential barrier of finite height hinders the symmetric exchange of a lone hydron between hydroxylic (proton-donating) and ketonic (protonaccepting) oxygen centers. This talk will discuss ongoing efforts to build upon the known structural and dynamical properties of tropolone so as to explore related multiple proton-transfer events that are mediated by successive formation and breaking of several hydrogen bonds. Of particular interest are weakly-bound complexes created in situ under “cold” molecular-beam conditions by docking amphoteric ligands (e.g., HF and HCOOH) into the reaction cleft of the TrOH substrate. Such species have the tantalizing possibility of undergoing double proton transfer,^{a} with resulting tunnelinginduced bifurcation of rovibronic features reflecting the intrinsic vibrational and/or electronic specificity of the attendant unimolecular transformation. Spectroscopic studies of several hydrogen-bound TrOH complexes through use of the richly structured Ã^{1}B_{2}−X̃^{1}A_{1}(π^{∗} ← π) absorption system will be presented, with complementary quantum-chemical calculations serving to guide the assignment and interpretation of observed spectral patterns.

aL. A. Burns, D. Murdock, and P. H. Vaccaro, Mol. Phys., 108, 1171 (2010).